16:
8(c3 - c) = c (6c + 1)

Answer by checkley75(3666)

8(c^3-c)=c(6c+1)
8c^3-8c=6c^2+c
8c^3-6c^2-9c=0
c(8c^2-6c-9)=0
c(4c+3)(2c-3)=0
c=0 answer
4c+3=0
4c=-3
c=-3/4 or -.75 answer
2c-3=0
2c=3
c=3/2 or 1.5 answer.